OpenStudy (yanasidlinskiy):

Solve and graph the inequality. 4/5x + 5 < -3

OpenStudy (yanasidlinskiy):

This is what I have done so far: \[\frac{ 4 }{ 5 }x +5 < -3\] I subtracted 5 from both sides and I'm left with \[\frac{ 4 }{ 5 }x < -8\] Now i don't know what to do next.

OpenStudy (yanasidlinskiy):

@mathmale

OpenStudy (mathmale):

Hello, Yana!! What I'd do here would be to multiply both left and right side of this inequality by (5/4). This will, in effect, "solve" the inequality for x. Try doing that, please.

OpenStudy (yanasidlinskiy):

I don't really get this. Would it be -15/-12??

OpenStudy (mathmale):

Please simplify this. Think: What is your goal here? You need to end up with an inequality involving x.

OpenStudy (mathmale):

\[\frac{ 4 }{ 5 }x < -8 \rightarrow (\frac{ 5 }{ 4 })\frac{ 4x }{ 5 }<(\frac{ 5 }{ 4 })(-8)\]

OpenStudy (yanasidlinskiy):

Oh...Ok...I got it now. I actually started doing that from the original problem. Would it be \[x \ge -10\]

OpenStudy (yanasidlinskiy):

No. It would be \[x < -10\] right??

OpenStudy (yanasidlinskiy):

@mathmale Am I right?

OpenStudy (whpalmer4):

\[\frac{4}{5}x < -8\]\[(\frac{5}{4})\frac{4}{5}x < (\frac{5}{4})(-8)\]\[\cancel{(\frac{5}{4})\frac{4}{5}}x < (\frac{-40}{4})\]\[x< -10\] Remember that if you multiply or divide an inequality by a negative number, you have to change the direction of the inequality sign. It didn't happen here, but it's important to remember that!